Vladimir opened an RRSP deposit account on December 1 2017, with a deposit of $852. He added $1,618 on July 1, 2019 and $2,883 on November 1, 2021. How much is in his account on January 31, 2025, if the deposit earns 6.82% compounded monthly
Answer #1

The monthly rate will be = 6.82/12 = 0.5683%. The future value equation will look like:

852 x 1.005683^86 + 1618 x 1.005683^67 + 2883 x 1.005683^39 = 7348.35